OpenStudy (anonymous):

i need help with this question Multiply (3.3 × 10^4) × (3 × 10^18). Express the answer in scientific notation. I have done the math, and here's what I have come up with 3.3 x 10^4 = 33,000 3 × 10^18 = 3000000000000000000 my answers are: 9.9 × 10^12 9.9 × 10^14 9.9 × 10^22 9.9 × 10^72 none of them seem to be right though...

OpenStudy (anonymous):

think of this way figure out what 3*3.3 is then since both numbers are being multiplied by 10^n, just add the two numbers. 10^4 and 10^18 is also 10^22. Makes sense?

OpenStudy (anonymous):

(3.3 * 3) * (10^4 * 10^18)

OpenStudy (anonymous):

yup... and when you muliply with exponents, you can add so simplify to: (9.9) * (10^22)

OpenStudy (anonymous):

so the answer would be 9.9 X 10^22?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

youre welcome. do you understand the process?

OpenStudy (anonymous):

Try to use this formula if we have a^m * a^n = a^(m+n) So 10^18 *10^4 = 10^(18+4)